# Federated Learning With Energy Harvesting Devices: An MDP Framework

### v1

@@ VIII Conclusion
In this paper, we have studied FL with EH devices. We have conducted convergence analysis to demonstrate the influence of transmission policy on the learning performance. Based on the convergence bound, we have formulated a joint device scheduling and power control problem to minimize the optimality gap under energy and bandwidth constraints. We have further modeled this optimization problem as an MDP and demonstrated that its optimal policy possesses a monotone structure. We have also proposed a low-complexity relax-and-truncate algorithm that is asymptotically optimal as the number of devices increases. Moreover, for unknown channels and harvested energy statistics, we have developed an SE-DQN algorithm that leverages the monotone structure of the optimal policy to improve the training performance. Finally, extensive numerical experiments have been presented to validate the theoretical results and the effectiveness of the proposed algorithms.
@@ VIII-A Proof of Theorem 1
According to the model update rule (11), we have
$\boldsymbol{w}_{t+1}=$ $\boldsymbol{w}_{t}-\eta\frac{\sum_{n=1}^{N}\beta_{n,t}\zeta_{n,t}{\boldsymbol{g}}_{n,t}}{\sum_{n=1}^{N}\beta_{n,t}\zeta_{n,t}}$ (57)
$=$ $\boldsymbol{w}_{t}-\eta\bigg{(}\frac{\sum_{n=1}^{N}\nabla F_{n}(\boldsymbol{w}_{t})}{N}$
$+\underbrace{\frac{\sum_{n=1}^{N}\beta_{n,t}\zeta_{n,t}{\boldsymbol{g}}_{n,t}}{\sum_{n=1}^{N}\beta_{n,t}\zeta_{n,t}}-\frac{\sum_{n=1}^{N}\nabla F_{n}(\boldsymbol{w}_{t})}{N}}_{\boldsymbol{\epsilon}_{t}}\bigg{)}$
$=$ $\boldsymbol{w}_{t}-\eta\left(\nabla F(\boldsymbol{w}_{t})+\boldsymbol{\epsilon}_{t}\right),$
where $\boldsymbol{\epsilon}_{t}$ denotes the model aggregation error induced by partial device participation and packet drops at time slot $t$.
According to Assumption 1, the loss gap between two adjacent time slots can be upper bounded by
$F(\boldsymbol{w}_{t+1})-F(\boldsymbol{w}_{t})$ (58)
$\leq$ $\nabla F(\boldsymbol{w}_{t})^{\mathsf{T}}(\boldsymbol{w}_{t+1}-\boldsymbol{w}_{t})+\frac{L}{2}\|\boldsymbol{w}_{t+1}-\boldsymbol{w}_{t}\|^{2}$
$=$ $-\eta\nabla F(\boldsymbol{w}_{t})^{\mathsf{T}}\left(\nabla F(\boldsymbol{w}_{t})+\boldsymbol{\epsilon}_{t}\right)+\frac{\eta^{2}L}{2}\|\nabla F(\boldsymbol{w}_{t})+\boldsymbol{\epsilon}_{t}\|^{2}$
$=$ $\left(\frac{\eta^{2}L}{2}-\eta\right)\left\|\nabla F(\boldsymbol{w}_{t})\right\|^{2}+\frac{\eta^{2}L}{2}\left\|\boldsymbol{\epsilon}_{t}\right\|^{2}$
$+\left(\eta^{2}L-\eta\right)\nabla F(\boldsymbol{w}_{t})^{\mathsf{T}}\boldsymbol{\epsilon}_{t}$
@@
$\mathrel{\stackrel{{\scriptstyle\makebox[0.0pt]{\mbox{\tiny(a)}}}}{{\leq}}}$ $-\frac{\eta}{2}\left\|\nabla F(\boldsymbol{w}_{t})\right\|^{2}+\frac{\eta}{2}\left\|\boldsymbol{\epsilon}_{t}\right\|^{2},$
where (a) is due to $-\nabla F(\boldsymbol{w}_{t})^{\mathsf{T}}\boldsymbol{\epsilon}_{t}\leq\frac{\left\|\nabla F(\boldsymbol{w}_{t})\right\|^{2}+\left\|\boldsymbol{\epsilon}_{t}\right\|^{2}}{2}$. According to Assumption 3, the $\ell_{2}$ norm of the model aggregation error is upper bounded by $\left\|\boldsymbol{\epsilon}_{t}\right\|^{2}\leq\frac{\rho}{N}\sum_{n=1}^{N}\frac{1-\beta_{n,t}\zeta_{n,t}}{D_{n}}$. Then, we take the expectation over (58) and we have
$\mathbb{E}\left[F(\boldsymbol{w}_{t+1})-F(\boldsymbol{w}_{t})\right]$ (59)
@@
$\leq$ $-\frac{\eta}{2}\left\|\nabla F(\boldsymbol{w}_{t})\right\|^{2}+\frac{\eta}{2}\mathbb{E}\left[\left\|\boldsymbol{\epsilon}_{t}\right\|^{2}\right]$
@@
$\leq$ $-\frac{\eta}{2}\left\|\nabla F(\boldsymbol{w}_{t})\right\|^{2}+\frac{\eta\rho}{2N}\sum_{n=1}^{N}\frac{\mathbb{E}\left[1-\beta_{n,t}\zeta_{n,t}\right]}{D_{n}}$
$=$ $-\frac{\eta}{2}\left\|\nabla F(\boldsymbol{w}_{t})\right\|^{2}+\frac{\eta\rho}{2N}\sum_{n=1}^{N}\frac{1-\beta_{n,t}(1-q_{n,t}(h_{n,t},p_{n,t}))}{D_{n}}.$
Based on the strong convexity in Assumption 2, we have that the Polyak-Lojasiewicz inequality holds [36], i.e.,
$\left\|\nabla F(\boldsymbol{w})\right\|^{2}\geq 2\mu\left(F(\boldsymbol{w})-F(\boldsymbol{w}^{*})\right),\forall\boldsymbol{w}\in\mathbb{R}^{d}.$ (60)
Combining (60) into (59), we can obtain
@@
$\mathbb{E}\left[F(\boldsymbol{w}_{t+1})-F(\boldsymbol{w}_{t})\right]$
$\leq$ $-\frac{\eta}{2}\left\|\nabla F(\boldsymbol{w}_{t})\right\|^{2}+\frac{\eta\rho}{2N}\sum_{n=1}^{N}\frac{1-\beta_{n,t}(1-q_{n,t}(h_{n,t},p_{n,t}))}{D_{n}}$
$\leq$ $-\eta\mu\left(F(\boldsymbol{w}_{t})-F(\boldsymbol{w}^{*})\right)$
$+\frac{\eta\rho}{2N}\sum_{n=1}^{N}\frac{1-\beta_{n,t}(1-q_{n,t}(h_{n,t},p_{n,t}))}{D_{n}}.$ (61)
By subtracting $F(\boldsymbol{w}^{*})$ on both sides of (VIII-A), we have
@@
$\mathbb{E}\left[F(\boldsymbol{w}_{t+1})-F(\boldsymbol{w}^{*})\right]$ (62)
$\leq$ $\left(1-\eta\mu\right)\left(F(\boldsymbol{w}_{t})-F(\boldsymbol{w}^{*})\right)$
$+\frac{\eta\rho}{2N}\sum_{n=1}^{N}\frac{1-\beta_{n,t}(1-q_{n,t}(h_{n,t},p_{n,t}))}{D_{n}}.$
Recursively applying (62) for $t\in\{1,\ldots,T\}$, we have
$\mathbb{E}\left[F(\boldsymbol{w}_{T+1}-F(\boldsymbol{w}^{*})\right]$ (63)
$\leq$ $A^{T}\left(F(\boldsymbol{w}_{1})-F(\boldsymbol{w}^{*})\right)$
$+\frac{\eta\rho}{2N}\sum_{t=1}^{T}A^{T-t}\sum_{n=1}^{N}\frac{1-\beta_{n,t}(1-q_{n,t}(h_{n,t},p_{n,t}))}{D_{n}},$
with $A=1-\eta\mu$.

### v2

@@ VIII Conclusion
In this paper, we have studied FL with EH devices. We conducted a convergence analysis to demonstrate the influence of transmission policy on the learning performance. Based on the convergence bound, we have formulated a joint device scheduling and power control problem to minimize the optimality gap under energy and bandwidth constraints. We have further modeled this optimization problem as an MDP and demonstrated that its optimal policy possesses a monotone structure. We have also proposed a low-complexity relax-and-truncate algorithm that is asymptotically optimal as the number of devices increases. Moreover, for unknown channels and harvested energy statistics, we have developed an SE-DQN algorithm that leverages the monotone structure of the optimal policy to improve the training performance. Finally, extensive numerical experiments have been presented to validate the theoretical results and the effectiveness of the proposed algorithms.
@@ VIII-A Proof of Theorem 1
According to Assumption 1, we have
$F_{n}(\bar{\boldsymbol{w}}_{t}^{(k+1)})$ (58)
@@
$\mathrel{\stackrel{{\scriptstyle\makebox[0.0pt]{\mbox{\tiny(a)}}}}{{\leq}}}$ $F_{n}({\boldsymbol{w}}_{n,t}^{(k)})+\bigl{\langle}\nabla F_{n}({\boldsymbol{w}}_{n,t}^{(k)}),\bar{\boldsymbol{w}}_{t}^{(k+1)}-{\boldsymbol{w}}_{n,t}^{(k)}\bigr{\rangle}$
$+\frac{L}{2}\bigl{\|}\bar{\boldsymbol{w}}_{t}^{(k+1)}-{\boldsymbol{w}}_{n,t}^{(k)}\bigr{\|}^{2}$
$\mathrel{\stackrel{{\scriptstyle\makebox[0.0pt]{\mbox{\tiny(b)}}}}{{\leq}}}$ $F_{n}({\boldsymbol{w}}^{*})+\bigl{\langle}\nabla F_{n}({\boldsymbol{w}}_{n,t}^{(k)}),\bar{\boldsymbol{w}}_{t}^{(k+1)}-{\boldsymbol{w}}^{*}\bigr{\rangle}$
$+\frac{L}{2}\bigl{\|}\bar{\boldsymbol{w}}_{t}^{(k+1)}-{\boldsymbol{w}}_{n,t}^{(k)}\bigr{\|}^{2}$
$\leq$ $F_{n}({\boldsymbol{w}}^{*})+\bigl{\langle}\nabla F_{n}({\boldsymbol{w}}_{n,t}^{(k)}),\bar{\boldsymbol{w}}_{t}^{(k+1)}-{\boldsymbol{w}}^{*}\bigr{\rangle}$
$+L\bigl{\|}\bar{\boldsymbol{w}}_{t}^{(k+1)}-\bar{\boldsymbol{w}}_{t}^{(k)}\bigr{\|}^{2}+L\bigl{\|}\bar{\boldsymbol{w}}_{t}^{(k)}-{\boldsymbol{w}}_{n,t}^{(k)}\bigr{\|}^{2},$
where (a) and (b) follow from the $L$-smoothness and convexity of the local loss functions, respectively. Then, according to the local model update rule (3), we have
$\frac{1}{N}\sum_{n=1}^{N}\left\langle\boldsymbol{g}_{n,t}^{(k)},\bar{\boldsymbol{w}}_{t}^{(k+1)}-{\boldsymbol{w}}^{*}\right\rangle$ (59)
$=$ $-\frac{1}{\eta}\left\langle\bar{\boldsymbol{w}}_{t}^{(k+1)}-\bar{\boldsymbol{w}}_{t}^{(k)},\bar{\boldsymbol{w}}_{t}^{(k+1)}-{\boldsymbol{w}}^{*}\right\rangle$
$=$ $\frac{1}{2\eta}\bigl{\|}\bar{\boldsymbol{w}}_{t}^{(k)}-{\boldsymbol{w}}^{*}\bigr{\|}^{2}-\frac{1}{2\eta}\bigl{\|}\bar{\boldsymbol{w}}_{t}^{(k+1)}-\bar{\boldsymbol{w}}_{t}^{(k)}\bigr{\|}^{2}$
$-\frac{1}{2\eta}\bigl{\|}\bar{\boldsymbol{w}}_{t}^{(k+1)}-{\boldsymbol{w}}^{*}\bigr{\|}^{2}.$
Combining (59) into (58), it follows that
$F(\bar{\boldsymbol{w}}_{t}^{(k+1)})-F({\boldsymbol{w}}^{*})$ (60)
$=$ $\frac{1}{N}\sum_{n=1}^{N}\left(F_{n}(\bar{\boldsymbol{w}}_{t}^{(k+1)})-F_{n}({\boldsymbol{w}}^{*})\right)$
$\leq$ $\frac{1}{N}\sum_{n=1}^{N}\bigl{\langle}\nabla F_{n}({\boldsymbol{w}}_{n,t}^{(k)}),\bar{\boldsymbol{w}}_{t}^{(k+1)}-{\boldsymbol{w}}^{*}\bigr{\rangle}$
$+L\bigl{\|}\bar{\boldsymbol{w}}_{t}^{(k+1)}-\bar{\boldsymbol{w}}_{t}^{(k)}\bigr{\|}^{2}+\frac{L}{N}\sum_{n=1}^{N}\bigl{\|}\bar{\boldsymbol{w}}_{t}^{(k)}-{\boldsymbol{w}}_{n,t}^{(k)}\bigr{\|}^{2}$
$\leq$ $\frac{1}{N}\sum_{n=1}^{N}\bigl{\langle}\nabla F_{n}({\boldsymbol{w}}_{n,t}^{(k)})-\boldsymbol{g}_{n,t}^{(k)},\bar{\boldsymbol{w}}_{t}^{(k+1)}-{\boldsymbol{w}}^{*}\bigr{\rangle}$
$+L\bigl{\|}\bar{\boldsymbol{w}}_{t}^{(k+1)}-\bar{\boldsymbol{w}}_{t}^{(k)}\bigr{\|}^{2}+\frac{L}{N}\sum_{n=1}^{N}\bigl{\|}\bar{\boldsymbol{w}}_{t}^{(k)}-{\boldsymbol{w}}_{n,t}^{(k)}\bigr{\|}^{2}$
$+\frac{1}{2\eta}\bigl{\|}\bar{\boldsymbol{w}}_{t}^{(k)}-{\boldsymbol{w}}^{*}\bigr{\|}^{2}-\frac{1}{2\eta}\bigl{\|}\bar{\boldsymbol{w}}_{t}^{(k+1)}-\bar{\boldsymbol{w}}_{t}^{(k)}\bigr{\|}^{2}$
$-\frac{1}{2\eta}\bigl{\|}\bar{\boldsymbol{w}}_{t}^{(k+1)}-{\boldsymbol{w}}^{*}\bigr{\|}^{2}.$
Moreover, given the unbiased stochastic gradient assumption in Assumption 2, i.e., $\mathbb{E}\left[\nabla F_{n}({\boldsymbol{w}}_{n,t}^{(k)})-\boldsymbol{g}_{n,t}^{(k)}\right]=0$, we have
@@
$\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}\left[\bigl{\langle}\nabla F_{n}({\boldsymbol{w}}_{n,t}^{(k)})-\boldsymbol{g}_{n,t}^{(k)},\bar{\boldsymbol{w}}_{t}^{(k+1)}-{\boldsymbol{w}}^{*}\bigr{\rangle}\right]$
@@
$=$ $\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}\left[\bigl{\langle}\nabla F_{n}({\boldsymbol{w}}_{n,t}^{(k)})-\boldsymbol{g}_{n,t}^{(k)},\bar{\boldsymbol{w}}_{t}^{(k+1)}-\bar{\boldsymbol{w}}_{t}^{(k)}\bigr{\rangle}\right]$
$\leq$ $\eta\mathbb{E}\left[\left\|\frac{1}{N}\sum_{n=1}^{N}\left(\nabla F_{n}({\boldsymbol{w}}_{n,t}^{(k)})-\boldsymbol{g}_{n,t}^{(k)}\right)\right\|^{2}\right]$
$+\frac{1}{4\eta}\mathbb{E}\left[\left\|\bar{\boldsymbol{w}}_{t}^{(k+1)}-\bar{\boldsymbol{w}}_{t}^{(k)}\right\|^{2}\right]$
$\leq$ $\frac{\eta\sigma_{l}^{2}}{N}+\frac{1}{4\eta}\mathbb{E}\left[\left\|\bar{\boldsymbol{w}}_{t}^{(k+1)}-\bar{\boldsymbol{w}}_{t}^{(k)}\right\|^{2}\right].$ (61)
Plugging (VIII-A) back to (60) yields
$\mathbb{E}\left[F(\bar{\boldsymbol{w}}_{t}^{(k+1)})-F({\boldsymbol{w}}^{*})\right]$
$\leq$ $\frac{\eta\sigma_{l}^{2}}{N}+\frac{1}{4\eta}\mathbb{E}\left[\left\|\bar{\boldsymbol{w}}_{t}^{(k+1)}-\bar{\boldsymbol{w}}_{t}^{(k)}\right\|^{2}\right]$
$+L\mathbb{E}\left[\bigl{\|}\bar{\boldsymbol{w}}_{t}^{(k+1)}-\bar{\boldsymbol{w}}_{t}^{(k)}\bigr{\|}^{2}\right]+\frac{L}{N}\sum_{n=1}^{N}\mathbb{E}\left[\bigl{\|}\bar{\boldsymbol{w}}_{t}^{(k)}-{\boldsymbol{w}}_{n,t}^{(k)}\bigr{\|}^{2}\right]$
$+\frac{1}{2\eta}\mathbb{E}\left[\bigl{\|}\bar{\boldsymbol{w}}_{t}^{(k)}-{\boldsymbol{w}}^{*}\bigr{\|}^{2}\right]-\frac{1}{2\eta}\mathbb{E}\left[\bigl{\|}\bar{\boldsymbol{w}}_{t}^{(k+1)}-\bar{\boldsymbol{w}}_{t}^{(k)}\bigr{\|}^{2}\right]$
$-\frac{1}{2\eta}\mathbb{E}\left[\bigl{\|}\bar{\boldsymbol{w}}_{t}^{(k+1)}-{\boldsymbol{w}}^{*}\bigr{\|}^{2}\right]$
$\leq$ $\frac{\eta\sigma_{l}^{2}}{N}+\left(L-\frac{1}{4\eta}\right)\mathbb{E}\left[\left\|\bar{\boldsymbol{w}}_{t}^{(k+1)}-\bar{\boldsymbol{w}}_{t}^{(k)}\right\|^{2}\right]$
$+\frac{L}{N}\sum_{n=1}^{N}\mathbb{E}\left[\bigl{\|}\bar{\boldsymbol{w}}_{t}^{(k)}-{\boldsymbol{w}}_{n,t}^{(k)}\bigr{\|}^{2}\right]$
$+\frac{1}{2\eta}\mathbb{E}\left[\bigl{\|}\bar{\boldsymbol{w}}_{t}^{(k)}-{\boldsymbol{w}}^{*}\bigr{\|}^{2}\right]-\frac{1}{2\eta}\mathbb{E}\left[\bigl{\|}\bar{\boldsymbol{w}}_{t}^{(k+1)}-{\boldsymbol{w}}^{*}\bigr{\|}^{2}\right]$
$\mathrel{\stackrel{{\scriptstyle\makebox[0.0pt]{\mbox{\tiny(a)}}}}{{\leq}}}$ $\frac{\eta\sigma_{l}^{2}}{N}+\frac{L}{N}\sum_{n=1}^{N}\mathbb{E}\left[\bigl{\|}\bar{\boldsymbol{w}}_{t}^{(k)}-{\boldsymbol{w}}_{n,t}^{(k)}\bigr{\|}^{2}\right]$
$+\!\frac{1}{2\eta}\mathbb{E}\left[\bigl{\|}\bar{\boldsymbol{w}}_{t}^{(k)}-{\boldsymbol{w}}^{*}\bigr{\|}^{2}\right]\!-\!\frac{1}{2\eta}\mathbb{E}\left[\bigl{\|}\bar{\boldsymbol{w}}_{t}^{(k+1)}-{\boldsymbol{w}}^{*}\bigr{\|}^{2}\right],$ (62)
where (a) holds due to $\eta\leq\frac{1}{4L}$. Summing the inequality (VIII-A) from $k=1$ to $K$, it follows that
$\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}\left[F(\bar{\boldsymbol{w}}_{t}^{(k+1)})-F({\boldsymbol{w}}^{*})\right]$
@@
$\leq$ $\frac{\eta\sigma_{l}^{2}}{N}+\frac{L}{NK}\sum_{n=1}^{N}\sum_{k=1}^{K}\mathbb{E}\left[\bigl{\|}\bar{\boldsymbol{w}}_{t}^{(k)}-{\boldsymbol{w}}_{n,t}^{(k)}\bigr{\|}^{2}\right]$
$+\frac{1}{2\eta K}\sum_{k=1}^{K}\left(\mathbb{E}\left[\bigl{\|}\bar{\boldsymbol{w}}_{t}^{(k)}-{\boldsymbol{w}}^{*}\bigr{\|}^{2}\right]-\mathbb{E}\left[\bigl{\|}\bar{\boldsymbol{w}}_{t}^{(k+1)}-{\boldsymbol{w}}^{*}\bigr{\|}^{2}\right]\right)$
@@
$\leq$ $\frac{\eta\sigma_{l}^{2}}{N}+\frac{L}{NK}\sum_{n=1}^{N}\sum_{k=1}^{K}\mathbb{E}\left[\bigl{\|}\bar{\boldsymbol{w}}_{t}^{(k)}-{\boldsymbol{w}}_{n,t}^{(k)}\bigr{\|}^{2}\right]$
$+\frac{1}{2\eta K}\left(\mathbb{E}\left[\bigl{\|}\bar{\boldsymbol{w}}_{t}^{(1)}-{\boldsymbol{w}}^{*}\bigr{\|}^{2}\right]-\mathbb{E}\left[\bigl{\|}\bar{\boldsymbol{w}}_{t}^{(K+1)}-{\boldsymbol{w}}^{*}\bigr{\|}^{2}\right]\right).$ (63)
Then, according to Lemma 2 in [38], by taking $\eta\leq\frac{1}{4L}$, we have
$\mathbb{E}\left[\bigl{\|}\bar{\boldsymbol{w}}_{t}^{(k)}-{\boldsymbol{w}}_{n,t}^{(k)}\bigr{\|}^{2}\right]\leq 18K^{2}\eta^{2}\sigma_{g}^{2}+4K\eta^{2}\sigma_{l}^{2}.$ (64)
By combining (64) with (VIII-A) and telescoping over $t=1,\dots,T$, we have
$\mathbb{E}\Biggl{[}\frac{1}{KT}\sum_{t=1}^{T}\sum_{k=1}^{K}\bigl{(}F(\bar{\boldsymbol{w}}_{t}^{(k)})-F(\boldsymbol{w}^{\star})\bigr{)}\Biggr{]}$
$\leq$ $\frac{\eta\sigma_{l}^{2}}{N}+4K\eta^{2}L\sigma_{l}^{2}+18K^{2}\eta^{2}L\sigma_{g}^{2}$
$+\frac{1}{2\eta KT}\sum_{t=1}^{T}\left(\mathbb{E}\!\left[\bigl{\|}\bar{\boldsymbol{w}}_{t}^{(1)}\!-\!{\boldsymbol{w}}^{*}\bigr{\|}^{2}\right]\!-\!\mathbb{E}\!\left[\bigl{\|}\bar{\boldsymbol{w}}_{t}^{(K+1)}\!-\!{\boldsymbol{w}}^{*}\bigr{\|}^{2}\right]\right)$
$=$ $\frac{\eta\sigma_{l}^{2}}{N}+4K\eta^{2}L\sigma_{l}^{2}+18K^{2}\eta^{2}L\sigma_{g}^{2}$
$+\frac{1}{2\eta KT}\!\sum_{t=1}^{T}\!\left(\mathbb{E}\!\left[\bigl{\|}\bar{\boldsymbol{w}}_{t}^{(1)}\!-\!{\boldsymbol{w}}^{*}\bigr{\|}^{2}\right]\!-\!\mathbb{E}\!\left[\bigl{\|}{\boldsymbol{w}}_{t+1}\!-\!{\boldsymbol{w}}^{*}\bigr{\|}^{2}\right]\right)$
$+\frac{1}{2\eta KT}\!\sum_{t=1}^{T}\!\left(\mathbb{E}\!\left[\bigl{\|}{\boldsymbol{w}}_{t+1}\!-\!{\boldsymbol{w}}^{*}\bigr{\|}^{2}\right]\!-\!\mathbb{E}\!\left[\bigl{\|}\bar{\boldsymbol{w}}_{t}^{(K+1)}\!-\!{\boldsymbol{w}}^{*}\bigr{\|}^{2}\right]\right)$
$\mathrel{\stackrel{{\scriptstyle\makebox[0.0pt]{\mbox{\tiny(a)}}}}{{\leq}}}$ $\frac{\eta\sigma_{l}^{2}}{N}+4K\eta^{2}L\sigma_{l}^{2}+18K^{2}\eta^{2}L\sigma_{g}^{2}$
$+\frac{\|\boldsymbol{w}_{1}-\boldsymbol{w}^{*}\|^{2}}{2\eta KT}-\frac{1}{2\eta KT}\mathbb{E}\left[\|\boldsymbol{w}_{T+1}-\boldsymbol{w}^{*}\|^{2}\right]$
$+\frac{G^{2}}{2\eta NKT}\sum_{t=1}^{T}\sum_{n=1}^{N}\left(1-\beta_{n,t}(1-q_{n,t}(h_{n,t},p_{n,t}))\right)$
$\leq$ $\frac{\|\boldsymbol{w}_{1}-\boldsymbol{w}^{*}\|^{2}}{2\eta KT}+\frac{\eta\sigma_{l}^{2}}{N}+4K\eta^{2}L\sigma_{l}^{2}+18K^{2}\eta^{2}L\sigma_{g}^{2}$
$+\frac{G^{2}}{2\eta NKT}\sum_{t=1}^{T}\sum_{n=1}^{N}\left(1-\beta_{n,t}(1-q_{n,t}(h_{n,t},p_{n,t}))\right),$
where (a) follows directly from the global model update rule defined in (11) and the corresponding broadcasting mechanism. Then, by taking $\eta=\frac{1}{4L\sqrt{KT}}$, we have
$\mathbb{E}\Biggl{[}\frac{1}{KT}\sum_{t=1}^{T}\sum_{k=1}^{K}\bigl{(}F(\bar{\boldsymbol{w}}_{t}^{(k)})-F(\boldsymbol{w}^{\star})\bigr{)}\Biggr{]}$ (66)
$\leq$ $\frac{2L\|\boldsymbol{w}_{1}-\boldsymbol{w}^{*}\|^{2}}{\sqrt{KT}}+\frac{\sigma_{l}^{2}}{4LN\sqrt{KT}}+\frac{\sigma_{l}^{2}}{4LT}+\frac{9\sigma_{g}^{2}K}{8LT}$
$+\frac{2LG^{2}}{N\sqrt{KT}}\sum_{t=1}^{T}\sum_{n=1}^{N}\left(1-\beta_{n,t}(1-q_{n,t}(h_{n,t},p_{n,t}))\right).$